OpenStudy (anonymous):

using this formula a^2+b^2=c^2 7 X 14 this is a triangle

OpenStudy (turingtest):

is there a question here somewhere?

OpenStudy (anonymous):

we need to find X

OpenStudy (turingtest):

is it a leg or the hypotenuse?

OpenStudy (turingtest):

if x is a leg than\[7^2+x^2=14^2\]if it is the hypotenuse then\[7^2+14^2=x^2\]either way solve for x

OpenStudy (anonymous):

if x is hypotenuse it is 15.65 if x is the leg x is 12.12

OpenStudy (anonymous):

it is the leg

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

leg = sqrt ( 12^2 - 7^2) = 9.75

OpenStudy (anonymous):

Thaknks